For the following electrochemical cell:

Co(s)/Co^2+(aq,0.0155M)//Ag^+(aq,2.50M)/Ag(s)

Write the net equation:______________

Calculate the following values at 25 degrees celcius using the standard potentials:

E standard cell ________V

E cell _____________V
Answer #1

Ag+ + e-\rightarrow Ag                Eo = +0.7996

Co \rightarrow Co2+ + e-              Eo = +0.28

----------------------

Ag+ + Co \rightarrow Ag + Co2+   Eo = 0.7996 + 0.28 = 1.0796 V = E standard cell

Q = [Co2+]/[Ag+] = (0.0155)/(2.50) = 6.2 \times 10-3

E = Eo - (0.0591/n)log Q

E = 1.0796 - (0.0591/2) log (6.2 \times 10-3) = 1.144 V = E cell